OpenStudy (anonymous):

Determine the solution of the system. Show your work. Y = 5x – 9 Y = 2x + 6

OpenStudy (acxbox22):

since both equations are equal to y we can set the equations equal 5x-9=2x+6 solve for x then sub the x value back intoeither equation and find y

OpenStudy (anonymous):

-9 + 5x + -2x = 6 + 2x + -2x